OpenStudy (anonymous):

What is the solution of the system? Use substitution. x = –2y + 10 –6y = 6x – 6 z = –3x + 10y A. x = 9, y = –8, z = 114 B. x = –8, y = 9, z = 114 C. x = –8, y = 114, z = 9 D. x = 114, y = 9, z = –8

OpenStudy (anonymous):

x = –2y + 10 –6y = 6x – 6 x is -2y + 10 By using the substitution method -6y = 6(-2y+10) -6 -6y = -12y+60-6 -6y = -12y + 54 -6y - -12y = 54 -6y + 12y = 54 6y = 54 y = 9 Then if y = 9 x = –2y + 10 x = -2(9)+10 x = -18 + 10 x = -8 Now you have the value of x and y, can find the z?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

no thats not it

OpenStudy (anonymous):

i think its B not C

OpenStudy (anonymous):

b. x = –8, y = 9, z = 114

OpenStudy (anonymous):

that was my next answer thanks

OpenStudy (anonymous):

your welcome